Solve the equation
2(x-3)-3x=4-5x

Answer by unlockmath(1688)

Hello,
2(x-3)-3x=4-5x
Expand:
2x-6-3x=4-5x
Combine:
-x-6=4-5x
Add 5x and 6 to get:
4x=10
x=10/4 or 2 1/2
Make sense?
RJ
www.math-unlock.com